If you mix ethanol with gasoline (C8H18) by 2:8 mass ratio, what is the A:F value ?
Answer #1

Given that ethanol is mixed with gasoline(C8H18) in the ratio 2:8.

So, lets take the basis as 10g, that is there is 2g of ethanol and 8g of gasoline,

Now we will calculate the required air amount using stoichiometric equation of ethanol and Gasoline reaction with Oxygen.

Сң, ені оң + 352 - cot + 2H2O

By this equation for 1 mol of ethanol it requires 3 moles of O2

So for 0.0435 mol ethanol we require 3*0.0435 mol=0.13 mol O2. -----(1)

We got 0.0435 by using the formula

no. of moles=given mass/molecular mass ie. (0.0435=2/46) -----eq(1)

Similarily for gasoline,

& His & - 8 CO2 & 9H20

1 mol gasoline requires 12.5 moles of O2

so 0.07 (8/114) mol of gasoline requires 12.5*0.07=0.877 moles O2   ----(2)

So we need a total of 0.13 + 0.877= 1 mol of O2

we know that air contains approximately 21% O2. So number of moles of air needed is 1/0.21 =4.762 moles air

Using the same eq 1 we can find the mass of air needed, by multiplying number of moles with its average molecular weight

that is 4.762*29 = 138.1 g of air.

so now air:fuel ratio of the given mixture is 138.1:10 or 13.:1